I am really having troubles with these word problems.
A small manufacturers weekly profit in dollars is given by P = 2x^2 + 200x. find the number of items of X that must be produced to make a profit of $4608

Answer by stanbon(75887)

A small manufacturers weekly profit in dollars is given by P = 2x^2 + 200x. find the number of items of X that must be produced to make a profit of $4608
-------
Solve:
2x^2 + 200x = 4608
-----
2x^2 + 200x - 4608 = 0
-----
x^2 + 100x - 2304 = 0
-----
x = [-100 +- sqrt(10000-4*-2304)]/2
----
x = [-100 +- sqrt(12304)]/2 = [-100 + 138.62]/2 = 38.62/2 = 19.31
-----
Cheers,
Stan H.
===============